# Seeing Sound, Hearing Sight: Uncovering Modality Bias and Conflict of AI models in Sound Localization

### v1

@@ 4 Discussion
We systematically and quantitatively examine modality biases and conflict resolution in sound source localization (SSL) across humans and AI models. Our study spans six audiovisual conditions, including congruent cues, conflicting audio-visual signals, and scenarios with missing audio or visual input in natural scenes. Our results reveal that object size significantly affects human performance. While conflicting or absent visual cues impair human SSL, people remain robust at successfully localizing even small sounding objects under challenging conditions, such as in multi-instance settings or with no visual input at all. In contrast, current multimodal AI models show strong dependence on visual cues. They often misattribute sound to large, centered, and audible objects and suffer sharp performance drops when visual cues are absent. Conflicting visual cues can even further degrade the model performance. Most models perform near chance when localizing small sound sources without accompanying visuals.
We hypothesize that these limitations stem from two key issues: the use of low-quality, biased audiovisual datasets and monaural audio inputs that lack spatial fidelity. To address this, we curate a high-quality dataset—AudioCOCO—by applying rigorous filtering criteria to audio and images. Using depth maps and 3D simulation engines, we generate realistic, spatialized stereo audio that obeys the physical laws of sound propagation. We fine-tune existing models on AudioCOCO using stereo inputs. Despite using fewer training examples, the resulting model surpasses current state-of-the-art across all conditions. Remarkably, without training on human behavior, the resulting model displays human-like localization biases, favoring horizontal over vertical precision. Ablation studies confirm the critical role of both dataset quality and stereo input in capturing spatial cues. This work highlights the importance of curating less biased datasets and building models that respect the constraints of physical sensory systems. Future directions include scaling up the dataset, incorporating temporal dynamics from video, and enhancing realism in simulated sound rendering.
@@
(a) The spatial distribution of single-object by category.
@@
Figure S1: The visualization results of images’ spatial distributions in the AudioCOCO dataset are presented for each object category, where the color bar indicates the frequency of object occurrences at each spatial location within the images. Darker regions correspond to higher frequencies.
@@
(b) The count distribution of single-object by category.
Figure S2: The visualization results of the image count distributions in the AudioCOCO dataset are presented, where blue, green, and red represent object size1, size2, and size3, respectively. This visualization illustrates how instances are distributed across different object size categories, highlighting the dataset’s balanced distribution in terms of object sizes.
More details on image selection. The COCO 2014 dataset offers annotations for 80 object categories, including both bounding boxes and segmentation masks. To ensure a fair comparison—especially given that some model backbones are pre-trained on COCO—we utilized the COCO 2014 validation set rather than the training set. From this, we selected 12 common vocal categories frequently encountered in daily life: person, motorbike, train, boat, elephant, bird, cat, dog, horse, sheep, cow, and keyboard. A total of 29,737 images containing at least one vocal object were extracted to form the pool of image candidates for our dataset.
@@
Additionally, given that the image distribution over all object categories in COCO is not uniform, we performed random sampling to cap the number of instances to 150 per category per object size bin, promoting balanced representation within the AudioCOCO dataset. The final distributions after filtering are shown in Fig. 2(a) and Fig. 2(b), demonstrating that AudioCOCO achieves a balanced dataset across object area sizes, object center positions, and categories.
More details on audio selection. To ensure that only high-quality, semantically aligned clips are retained, we apply the following three filtering steps to the videos from VGGSound vggsound . First, we introduce Semantic Consistency (SeC) and select audios that are representative of the semantic object categories. Specifically, we take all audio files belonging to the same semantic category, extract audio features from the last layer of the Wav2vec baevski2020wav2vec model, compute their pari-wise cosine similarities based on these features, and retain the top 80% audios with the highest cosine similarity scores. Second, to ensure the audio is free from noise or interference from unrelated sources, we introduce Mel-Spectrogram Similarity (MSS) as a filtering criterion within each category. For each audio clip, we compute its Mel spectrogram, average the frequency magnitudes over time, and apply a logarithmic transformation to compress high-frequency components—yielding a compact representation of the audio’s overall spectral structure. We then calculate the cosine similarity between these representations and retain the top 65% of audio clips based on this MSS metric, following the SeC criteria. Third, to ensure the stereo audio corresponds to the sounding object being centered in the video frame, we introduce Spatial Consistency (SpC) for each audio-image pair. We compute the Spearman correlation between the left and right audio channels; a high correlation suggests the sound source is centrally located, producing similar waveforms in both channels. We retain the top 50% of audio clips based on this metric from the MSS-filtered set.
@@
(a) The three stages of audio filtering result.
(b) The comparison between original audio counts and filtering audio counts
Figure S3: The visualization results of audio statistics in the AudioCOCO dataset demonstrate that the audio quality across most categories has significantly improved following the filtering process. Additionally, we ensured a balanced count distribution among all categories. These outcomes highlight the effectiveness of our selection and refinement strategy in enhancing both the semantic consistency and acoustic quality of the dataset.
@@
SSLTI ssltie , LVS vggss , FNAC fnac , and IS3 is3 are SSL models based on dual-stream architectures with 2D Convolutional Neural Networks (CNNs). Each model processes visual and auditory inputs separately using dedicated encoders before fusing the features for contrastive learning during training. Both encoders are based on ResNet18 resnet . Beyond standard contrastive learning, IS3 introduces an Intersection-over-Union (IoU) loss and a semantic alignment loss to improve localization accuracy and better alignment between audio and visual modalities during supervised training. All the models are trained on the FlickrSoundNet aytar2016soundnet and VGG-Sound vggsound datasets.
CAVP cavp is a sound source segmentation model that follows a dual-stream CNN architecture. It uses PVTV2-B5 wang2022pvt as the visual encoder and either VGGish vggish or ResNet18 as the audio encoder. CAVP is trained in a fully supervised manner using cross-entropy loss for segmentation and contrastive learning to align audio-visual features. Training data includes AVSBench zhou2024audio and VPO cavp datasets.
AVSegformer gao2024avsegformer is an audio-visual semantic segmentation model built on a dual-stream transformer-based architecture. It employs SegFormer xie2021segformer as the visual backbone and Audio Spectrogram Transformer (AST) ast as the audio encoder to capture fine-grained cross-modal representations. AVSegformer integrates both modality-specific and fused token embeddings through a lightweight fusion decoder for pixel-level prediction. It is trained in a fully supervised setting using a combination of cross-entropy loss for segmentation and audio-visual consistency loss. AVSegformer’s training data includes three subsets of AVSBench.
ImageBind imagebind and LanguageBind languagebind are large-scale, multi-modal transformer models with billions of parameters. These models are trained with contrastive learning objectives across six modalities: image, audio, video, text, depth, thermal, and IMU, and embed these into a shared representation space. Both models use a Vision Transformer(ViT) vit for visual encoding and AST ast for audio encoding. ImageBind is trained on large-scale datasets including LAION laion ; cherti2023reproducible , SSv2 ssv2 , and K400 k400 . LanguageBind builds on the pre-trained encoders from ImageBind and further fine-tunes them on the VIDAL-10M languagebind dataset.
@@
Figure S6: The figure shows the failure examples of both IS3 and IS3-FT-Stereo compared to humans. The leftmost images (Columns 1 and 4) show the correct localization results made by humans. Red boxes mark the ground truth sound source locations, while green circles indicate mouse click responses from multiple participants. The middle columns (2 and 5) and rightmost columns (3 and 6) display heatmaps predicted by IS3-FT-Stereo and IS3, respectively. Yellow triangles on the heatmaps denote the predicted sound source locations from each model. See the colorbar for the activation values of the heatmaps.
Implementation details of AI models. All ten models—except IS3-FT-Stereo—take paired image and mono audio inputs, processing each modality independently through dedicated visual and audio encoders. For the mono audio setup, we follow the IS3-FT-Mono design by averaging the stereo channels into a single-channel input. For CNN-based models, feature maps from the final layers of the encoders are extracted, and cosine similarity is computed between them to produce a similarity heatmap. For transformer-based models, token sequences from both encoders are retrieved, and pairwise dot products of their token embeddings are calculated to generate the heatmap. The predicted sound source location is identified as the point with the highest activation on the heatmap.
We evaluated all the current models using their publicly available, pre-trained weights and adhered to their original implementation details. Since AVSBench zhou2024audio shares some images with our AudioCOCO test set, we exclude those overlapping image-audio pairs when evaluating CAVP cavp and AVSegformer gao2024avsegformer . Experiments for IS3-FT, ImageBind and LanguageBind were conducted on 8 NVIDIA A100 GPUs, while all other models were tested on 4 NVIDIA A6000 GPUs. To fine-tune IS3-FT models, we use the SGD optimizer with a momentum of 0.9 and weight decay of 0.0005. The initial learning rate is set to 0.001, and the batch size is 32. All the model evaluations are run 3 times with different random seeds.
@@
Congruent Conflicting VCue Absent VCue Audio Only Vision Only Multi-Instance Model Acc(%) A-Acc Vision Gray Vision Noise Audio Silent Audio Noise A-Acc V-Acc Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Random 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.5 21.9 1.6 9.1 21.3 8.4 17.8 26.2 IS3is3 4.3 13.5 31.0 4.3 7.9 14.5 6.2 11.0 20.8 1.4 5.2 9.4 1.5 5.4 10.7 25.9 32.0 69.5 30.3 58.0 77.6 4.2 10.3 22.4 17.2 22.8 49.5 FNACfnac 3.8 11.5 26.9 3.9 7.2 13.3 5.8 10.5 18.6 1.1 4.8 8.7 1.3 5.0 9.6 23.2 26.5 63.8 27.4 52.6 71.9 0.7 3.9 26.3 13.4 25.5 52.8 SSLTIEssltie 3.4 9.9 23.7 3.1 6.8 12.4 5.2 9.9 16.5 0.7 3.9 8.2 1.1 4.3 8.9 23.1 26.5 60.7 25.5 49.7 69.8 0.6 2.5 24.0 13.2 20.7 49.4 LVSvggss 3.3 9.7 23.6 3.0 6.5 12.1 5.0 9.4 16.0 0.7 3.8 8.3 1.0 4.2 8.8 23.2 26.4 60.5 25.4 49.7 70.2 0.5 3.9 25.3 12.5 22.4 50.6 CAVPcavp 5.1 15.2 38.7 3.8 7.0 13.2 5.5 10.2 18.4 0.9 4.7 8.4 1.2 4.9 9.8 25.4 30.7 70.5 29.4 56.5 72.3 5.9 11.5 27.4 20.5 26.0 59.5 AVSegformergao2024avsegformer 4.7 15.5 39.5 3.6 7.1 12.8 5.3 9.8 17.2 0.9 4.5 7.9 1.1 4.6 9.4 25.0 30.3 69.9 28.7 55.8 70.6 5.5 11.3 26.5 19.6 24.7 56.3 ImageBindimagebind 1.4 5.8 14.5 2.2 5.7 11.3 3.4 7.2 12.8 0.3 3.1 6.6 0.4 3.8 7.5 19.8 24.5 47.2 21.2 26.3 52.0 0.5 2.7 15.3 6.5 12.4 25.1 LanguageBindlanguagebind 1.1 4.6 12.7 2.1 5.4 10.7 3.2 6.9 10.6 0.2 3.0 6.5 0.4 3.7 7.6 18.6 23.2 44.9 19.1 23.8 48.8 0.4 2.1 13.0 6.2 10.8 22.5   IS3-FT -Stereo 19.3 29.1 52.2 8.8 14.7 27.5 10.2 18.2 31.0 3.1 12.3 23.7 3.8 13.5 26.0 40.6 61.1 83.2 45.5 76.8 90.5 8.9 14.3 31.5 85.4 89.5 93.7 Human 32.9 56.6 69.3 22.8 39.5 47.1 12.7 35.3 41.2 4.5 30.3 45.6 4.2 28.8 44.9 - - 18.2 30.5 52.8 97.5 99.1 99.8 Table S1: Comparison of AI models in UniAV. Bold and underlined values indicate the best and second-best performances, respectively.
As illustrated in Fig. S6, we further highlight the limitations of IS3-FT-Stereo in comparison to human performance. Specifically, IS3-FT-Stereo struggles to localize the correct sounding object when a distractor occupies a large visual region in the scene. This suggests that the model is heavily influenced by visual saliency, often prioritizing visually prominent objects over auditory cues, a bias that humans are better able to overcome.

### v2

@@ 4 Discussion
We systematically and quantitatively examine modality biases and conflicts in SSL across humans and AI models. Our study covers six audiovisual conditions, including congruent cues, conflicting signals, and cases with missing audio or visual input in natural scenes. Human listeners show strong robustness: although conflicting or absent visual cues reduce performance, they can still accurately localize even small sound sources under challenging or multi-instance conditions, and even in the absence of visual input. In contrast, current multimodal AI models rely heavily on vision—misattributing sounds to large, centered, and salient objects, and suffering steep performance drops when visual cues are removed. Conflicting visuals further degrade their accuracy, with most models performing near chance for small or visually absent sound sources.
We identify two primary causes of these limitations: (1) low-quality, visually biased audiovisual datasets, and (2) monaural audio inputs lacking spatial fidelity. To overcome these issues, we curate AudioCOCO, a high-quality dataset built through rigorous filtering and 3D physical simulation. By integrating depth maps, HRTF-based filtering that mimics pinna effects, cochlear-inspired frequency decomposition and modulation, and physically grounded 3D sound propagation, AudioCOCO produces realistic, spatialized stereo audio aligned with human auditory processing. Building on this, we introduce EchoPin, a neuroscience-inspired model trained on AudioCOCO. Despite fewer training examples, EchoPin surpasses state-of-the-art models across all conditions and exhibits human-like localization biases, such as stronger precision along the horizontal plane. Ablation studies confirm the importance of both high-quality datasets and stereo auditory input for capturing spatial cues. This work underscores the value of designing models and datasets that respect the physical constraints of sensory systems. Future directions include scaling AudioCOCO to incorporate temporal dynamics from videos and improving realism in simulated sound rendering, such as the effect of refraction.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its NRFF award NRF-NRFF15-2023-0001 and Mengmi Zhang’s Startup Grant from Nanyang Technological University, Singapore. We would also like to thank Qing Lin and Shuangpeng Han for their valuable advice and feedback on the project.
Supplementary Material
In the supplementary material, we provide additional information of our AudioCOCO dataset in Sec.S1 and human experiment setup in Sec.S2. Moreover, we include extra experimental results, an ablation study, and visualization results for predicted position bias, which are discussed in Sec.S4 and Sec.S5.
@@ Acknowledgements
(a) The spatial distribution of single objects by category.
@@
Figure S1: The visualization results of images’ spatial distributions in the AudioCOCO test set are presented for each object category, where the color bar indicates the frequency of object occurrences at each spatial location within the images. Darker regions correspond to higher frequencies.
@@
(b) The count distribution of single objects by category.
Figure S2: The visualization results of the image count distributions in the AudioCOCO test set are presented, where blue, green, and red represent object size1, size2, and size3, respectively. This visualization illustrates how instances are distributed across different object size categories, highlighting the dataset’s balanced distribution in terms of object sizes.
(a) The spatial distribution of single objects by category for AudioCOCO’s training set.
(b) The count distribution of single objects by category for AudioCOCO’s training set.
Figure S3: The visualization results of the image count distribution and spatial distribution for the AudioCOCO training set. For the count distribution, object sizes are color-coded: blue for Size1, green for Size2, and red for Size3. This plot illustrates the number of object instances across different size categories, confirming that the dataset maintains a balanced distribution in terms of object size. For the spatial distribution, the heatmap reflects the frequency of object occurrences at each spatial location in the images, with the color bar indicating frequency. Darker regions correspond to areas of higher object density, highlighting positional biases or spread across the dataset.
More details on image selection. The MSCOCO 2014 dataset coco offers annotations for 80 object categories, including both bounding boxes and segmentation masks. To ensure a fair comparison—especially given that some model backbones are pre-trained on MSCOCO—we utilized the MSCOCO 2014 validation set rather than the training set. From this, we selected 12 common vocal categories frequently encountered in daily life: person, motorbike, train, boat, elephant, bird, cat, dog, horse, sheep, cow, and keyboard. A total of 29,737 images containing at least one vocal object were extracted to form the pool of image candidates for our dataset.
@@
Additionally, given that the image distribution over all object categories in COCO is not uniform, we performed random sampling to cap the number of instances to 150 per category per object size bin, promoting balanced representation within the AudioCOCO dataset. The final distributions after filtering are shown in Fig. 2(a), Fig. 2(b), and Fig. S3, demonstrating that AudioCOCO achieves a balanced dataset across object area sizes, object center positions, and categories.
More details on audio selection. To ensure that only high-quality, semantically aligned clips are retained, we apply the following three filtering steps to the videos from VGGSound (vggsound, ). First, we introduce Semantic Consistency (SeC) and select audios that are representative of the semantic object categories. Specifically, we take all audio files belonging to the same semantic category, extract audio features from the last layer of the Wav2vec (baevski2020wav2vec, ) model, compute their pari-wise cosine similarities based on these features, and retain the top 80% audios with the highest cosine similarity scores. Second, to ensure the audio is free from noise or interference from unrelated sources, we introduce Mel-Spectrogram Similarity (MSS) as a filtering criterion within each category. For each audio clip, we compute its Mel spectrogram, average the frequency magnitudes over time, and apply a logarithmic transformation to compress high-frequency components—yielding a compact representation of the audio’s overall spectral structure. We then calculate the cosine similarity between these representations and retain the top 65% of audio clips based on this MSS metric, following the SeC criteria. Third, to ensure the stereo audio corresponds to the sounding object being centered in the video frame, we introduce Spatial Consistency (SpC) for each audio-image pair. We compute the Spearman correlation between the left and right audio channels; a high correlation suggests the sound source is centrally located, producing similar waveforms in both channels. We retain the top 50% of audio clips based on this metric from the MSS-filtered set.
@@
(a) The three stages of audio filtering result for AudioCOCO test set.
(b) The comparison between original audio counts and filtering audio counts for AudioCOCO test set.
Figure S4: The visualization results of audio statistics in the AudioCOCO test set demonstrate that the audio quality across most categories has significantly improved following the filtering process. Additionally, we ensured a balanced count distribution among all categories. These outcomes highlight the effectiveness of our selection and refinement strategy in enhancing both the semantic consistency and acoustic quality of the dataset.
(a) The three stages of audio filtering result for AudioCOCO training set.
(b) The comparison between original audio counts and filtering audio counts for AudioCOCO training set.
Figure S5: The visualization results of audio statistics in the AudioCOCO training set demonstrate that the audio quality across most categories has significantly improved following the filtering process. Additionally, we ensured a balanced count distribution among all categories. These outcomes highlight the effectiveness of our selection and refinement strategy in enhancing both the semantic consistency and acoustic quality of the dataset.
@@
Center fixation presentation before the visual stimulus onset. During the experiment, participants were instructed to fixate on a central dot before each trial began. This pre-trial fixation is a standard element in human psychophysics and cognitive neuroscience, designed to recenter attention and minimize carry-over effects across trials. By requiring participants to begin each trial from a common spatial and attentional baseline, this design ensures that any differences in response latency or eye movement patterns can be attributed to the experimental manipulation, rather than lingering attentional bias from the previous trial.
@@
Figure S7: Overview of SOTA multi-modal models for sound source localization (SSL). All models receive paired images and mono audio inputs, where stereo signals are averaged into a single channel. Visual and auditory inputs are processed independently through separate encoders. For CNN-based models (indicated by blue arrows), feature maps from the final layers of each encoder are extracted, compared via cosine similarity, and used to generate a similarity heatmap. For transformer-based models, output token sequences are obtained from both encoders. Dot products on their token embeddings are calculated, and a heatmap is produced accordingly. During evaluation, SSL accuracy is determined by verifying whether the location of the maximum activation on the heatmaps (red triangles or red tokens) lies within the segmentation mask of the ground truth sounding object (red bounding box).
@@
SSLTI (ssltie, ), LVS (vggss, ), FNAC (fnac, ), and IS3 (is3, ) are SSL models based on dual-stream architectures with 2D Convolutional Neural Networks (CNNs). Each model processes visual and auditory inputs separately using dedicated encoders before fusing the features for contrastive learning during training. Both encoders are based on ResNet18 (resnet, ). Beyond standard contrastive learning, IS3 introduces an Intersection-over-Union (IoU) loss and a semantic alignment loss to improve localization accuracy and better alignment between audio and visual modalities during supervised training. All the models are trained on the FlickrSoundNet (aytar2016soundnet, ) and VGG-Sound (vggsound, ) datasets.
CAVP (cavp, ) is a sound source segmentation model that follows a dual-stream CNN architecture. It uses PVTV2-B5 (wang2022pvt, ) as the visual encoder and either VGGish (vggish, ) or ResNet18 as the audio encoder. CAVP is trained in a fully supervised manner using cross-entropy loss for segmentation and contrastive learning to align audio-visual features. Training data includes AVSBench (zhou2024audio, ) and VPO (cavp, ) datasets.
AVSegformer (gao2024avsegformer, ) is an audio-visual semantic segmentation model built on a dual-stream transformer-based architecture. It employs SegFormer (xie2021segformer, ) as the visual backbone and Audio Spectrogram Transformer (AST) (ast, ) as the audio encoder to capture fine-grained cross-modal representations. AVSegformer integrates both modality-specific and fused token embeddings through a lightweight fusion decoder for pixel-level prediction. It is trained in a fully supervised setting using a combination of cross-entropy loss for segmentation and audio-visual consistency loss. AVSegformer’s training data includes three subsets of AVSBench.
ImageBind (imagebind, ) and LanguageBind (languagebind, ) are large-scale, multi-modal transformer models with billions of parameters. These models are trained with contrastive learning objectives across six modalities: image, audio, video, text, depth, thermal, and IMU, and embed these into a shared representation space. Both models use a Vision Transformer(ViT) (vit, ) for visual encoding and AST (ast, ) for audio encoding. ImageBind is trained on large-scale datasets including LAION (laion, ; cherti2023reproducible, ), SSv2 (ssv2, ), and K400 (k400, ). LanguageBind builds on the pre-trained encoders from ImageBind and further fine-tunes them on the VIDAL-10M (languagebind, ) dataset.
@@
Implementation details of AI models. All eleven models—except EchoPin-S and EchoPin—take paired image and mono audio inputs, processing each modality independently through dedicated visual and audio encoders. For the mono audio setup, we follow the EchoPin-M design by averaging the stereo channels into a single-channel input. As shown in Fig. S7, for CNN-based models, feature maps from the final layers of the encoders are extracted, and cosine similarity is computed between them to produce a similarity heatmap. For transformer-based models, token sequences from both encoders are retrieved, and pairwise dot products of their token embeddings are calculated to generate the heatmap. The predicted sound source location is identified as the point with the highest activation on the heatmap.
We evaluated all the current models using their publicly available, pre-trained weights and adhered to their original implementation details. Since AVSBench (zhou2024audio, ) shares some images with our AudioCOCO test set, we exclude those overlapping image-audio pairs when evaluating CAVP (cavp, ) and AVSegformer (gao2024avsegformer, ). Experiments for ImageBind and LanguageBind were conducted on 8 NVIDIA A100 GPUs, whereas all other models, including EchoPin and its variants, were trained and evaluated on 4 NVIDIA A6000 GPUs. We fine-tune the EchoPin models using the Adaptive Moment Estimation (Adam) optimizer kinga2015method with a weight decay of $1\times 10^{-4}$ for 10 epochs. The initial learning rate is set to $1\times 10^{-5}$, and the batch size is 16. Each fine-tuning session takes approximately 16 hours to complete. To accelerate data loading, all audio waveforms are preprocessed and stored as cochleagram tensors in advance, with each .npy file occupying roughly 160 MB. All model evaluations are repeated three times with different random seeds to ensure statistical reliability.
@@
As shown in Fig.S9 and Fig.S10, we further illustrate the limitations of EchoPin relative to human performance and provide additional qualitative results for other baselines. Notably, EchoPin often fails to localize the correct sounding object when a visually salient distractor occupies a large portion of the scene. This suggests that the model remains vulnerable to visual saliency bias, tending to prioritize large or central objects even when they are not the true auditory source—a tendency that human listeners are better equipped to suppress.
Similar patterns are shown in CAVP, which exhibits modality conflict sensitivity akin to IS3. Both models are frequently misled by conflicting audio-visual cues and demonstrate a systematic bias toward the visual modality, highlighting a lack of robust auditory grounding under incongruent conditions.
Figure S9: The figure shows the failure examples of both IS3 and EchoPin compared to humans. The leftmost images (Columns 1 and 4) show the correct localization results made by humans. Red boxes mark the ground truth sound source locations, while green circles indicate mouse click responses from multiple participants. The middle columns (2 and 5) and rightmost columns (3 and 6) display heatmaps predicted by EchoPin and IS3, respectively. Yellow triangles on the heatmaps denote the predicted sound source locations from each model. See the colorbar for the activation values of the heatmaps.
Figure S10: The figure shows the failure examples of CAVP and EchoPin compared to humans. The rightmost columns (3 and 6) display heatmaps predicted by CAVP.
Appendix S5 More quantitative results of AI models in SSL
The raw A-Acc may be confounded by the area of the ground-truth mask—since, intuitively, smaller objects are more difficult to localize while larger ones are easier. To address this potential bias, we introduce a chance-corrected gain metric, which quantifies the improvement of a human or model over a random guess, normalized by the baseline accuracy of the random guess:
$\mathrm{Gain}(X)=\frac{\mathrm{Acc}_{X}-\mathrm{Acc}_{\text{rand}}}{\mathrm{Acc}_{\text{rand}}}\times 100$ (1)
The $\mathrm{Acc}_{\text{rand}}$ represents the accuracy achieved when predictions are sampled uniformly at random across the scene.
Normalized Percentage Size 1 Size 2 Size 3
(EchoPin - Random) / Random 5.2 4.6 2.9
(Human - Random) / Random 31.9 7.3 3.6
(Human - EchoPin) / Random 26.7 3.0 0.8
Table S1: We report the chance-corrected gain for humans and models across three object sizes under the congruent condition. While human performance increases significantly with larger object sizes, models show limited improvement. This discrepancy may stem from AudioCOCO’s uniform object size distribution, which limits the models’ ability to exploit size-dependent cues.
This normalization removes the influence of mask size and isolates the true localization capability of humans and models. As shown in Tab.S1, even after correcting for chance, humans consistently outperform models—especially for small object sizes. For example, in the smallest size category, humans achieve a gain of 31.9%, compared to only 5.2% for models. These results suggest that object size modulates sound source localization performance in a way that cannot be fully explained by ground-truth mask area alone, highlighting deeper perceptual and representational differences between human and model behavior.
We apply the same evaluation criterion to both human participants and AI models to ensure fair comparisons: a prediction is considered correct if the peak activation (for models) or the human click falls within the bounding box of the sounding object. We also conducted an additional analysis by varying the pixel distance thresholds used to determine correctness. Specifically, a prediction is considered correct if it falls within $x$ pixels of the ground-truth bounding box. We report the resulting A-Acc (accuracy with spatial tolerance) as a function of pixel thresholds in Tab.S2 based on congruent conditions for object size2. From these results, we observe that while larger thresholds naturally lead to higher A-Acc values, the relative performance trend between humans and models remains consistent. This further supports the validity of our evaluation methodology.
Threshold Random Human EchoPin-S EchoPin
0 (default) 9.5 79.3 17.3 50.9
10 9.6 79.9 18.5 52.0
25 9.9 80.4 19.8 52.4
Table S2: We compare performance across varying localization thresholds (0–25 pixels) under the congruent condition for object size 2. As shown, both human and model accuracy exhibit only marginal improvement with increasing thresholds, indicating that our evaluation metric is stable and not overly sensitive to small shifts in the decision boundary—thereby validating its robustness.
Method VGG-$S$ Flickr-SoundNet
IS3 42.96 84.40
CAVP 43.58 85.03
EchoPin-S 43.61 85.25
EchoPin 45.02 85.87
Table S3: We conduct comparison experiments on the VGG-SS and Flickr-SoundNet datasets. Despite the challenges posed by these large-scale, imbalanced public benchmarks, EchoPin consistently outperforms prior state-of-the-art (SOTA) methods, highlighting the robustness of spatial cues and the effectiveness of our neuroscience-inspired design.
Moreover, we conducted an additional experiment by fine-tuning our EchoPin-S model using the VGG-SS and Flickr-SoundNet datasets, and report comparative results against IS3 and CAVP using the CIoU metric under the default evaluation protocol provided by (avl2, ). As a localization-aware metric, CIoU accounts for overlap area, center distance, and aspect ratio alignment, offering a more comprehensive assessment of spatial prediction quality. From Tab. S3, we observe that benefiting from the fine-grained spatial features provided by AudioCOCO, our EchoPin-S model achieves superior performance on these standard SSL benchmarks, outperforming the baselines.
@@
Congruent Conflicting VCue Absent VCue Audio Only Vision Only Multi-Instance Model Acc(%) A-Acc Vision Gray Vision Noise Audio Silent Audio Noise A-Acc V-Acc Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Size1 Size2 Size3 Random 1.8 9.5 19.6 1.8 9.5 19.6 1.8 9.5 19.6 1.8 9.5 19.6 1.8 9.5 19.6 1.8 9.5 19.6 1.8 9.5 19.6 1.6 9.1 21.3 8.4 17.8 26.2 IS3 2.3 16.5 31.1 1.7 12.3 24.8 3.3 13.5 29.8 2.3 8.8 11.9 3.3 17.3 39.7 21.5 35.2 42.7 20.5 31.6 38.9 4.8 7.9 22.4 11.9 24.1 40.9 FNAC 1.8 12.5 26.9 1.9 7.2 13.3 5.8 10.5 18.6 1.1 4.8 8.7 1.3 5.0 9.6 6.2 16.5 43.8 5.4 12.6 11.9 0.7 3.9 26.3 13.4 25.5 52.8 SSLTIE 1.4 9.9 23.7 2.1 6.8 12.4 5.2 9.9 16.5 0.7 3.9 8.2 1.1 4.3 8.9 6.1 16.5 40.7 5.5 12.7 12.8 0.6 2.5 24.0 13.2 20.7 49.4 LVS 1.3 9.7 23.6 2.0 6.5 12.1 5.0 9.4 16.0 0.7 3.8 8.3 1.0 4.2 8.8 6.2 16.4 40.5 5.4 12.7 12.2 0.5 3.9 25.3 12.5 22.4 50.6 CAVP 3.1 15.2 38.7 3.8 7.0 13.2 5.5 10.2 18.4 0.9 4.7 8.4 1.2 4.9 9.8 8.4 21.2 44.5 6.6 16.5 36.3 2.9 7.5 20.4 10.5 23.0 39.5 AVSegformer 2.7 15.5 39.5 3.6 7.1 12.8 5.3 9.8 17.2 0.9 4.5 7.9 1.1 4.6 9.4 8.0 20.3 29.9 8.7 17.8 32.6 2.5 7.3 20.2 11.2 23.7 40.9 ImageBind 1.4 5.8 14.5 1.2 5.7 11.3 3.4 7.2 12.8 0.3 3.1 6.6 0.4 3.8 7.5 5.8 14.5 27.2 4.2 16.3 24.0 0.5 2.7 15.3 6.5 12.4 25.1 LanguageBind 1.1 4.6 12.7 1.1 5.4 10.7 3.2 6.9 10.6 0.2 3.0 6.5 0.4 3.7 7.6 5.6 13.2 24.9 4.1 13.8 24.8 0.4 2.1 13.0 6.2 10.8 22.5   EchoPin 11.2 50.9 76.0 16.6 21.9 26.4 11.8 54.0 77.2 10.0 16.9 17.9 4.0 20.2 40.8 26.3 39.0 53.7 13.5 17.6 21.9 4.5 24.1 47.1 37.5 53.8 64.2 Human 59.3 79.3 90.7 40.0 57.9 67.9 54.3 65.7 80.0 9.3 49.4 76.8 12.3 45.1 80.3 - - 25.7 36.4 38.6 60.9 82.8 89.1 Table S4: Comparison of AI models in our benchmark across six conditions and three object sizes. Bold and underlined values indicate the best and second-best performances, respectively. Since CAVP and AVSegformer are trained on MS-COCO2017, we evaluate them on a subset of AudioCOCO and denote their results with the symbol .